OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, 4) and is parallel to y = one thirdx - 3. y = 3x - 14 y = one thirdx + 2 y = -3x + 22 y = -one thirdx + 8 Can someone please help me. im really confused

OpenStudy (shaik0124):

just a min

OpenStudy (shaik0124):

Two different lines are parallel if they have the same slope. The slope of the line y = -3x - 4 is -3. Using the point-slope form of the equation of a line gives y - 6 = -3(x - 2) . The slope-intercept form is y = -3x + 12. Another way to solve the problem is let y = -3x + b and determine b from the given point.

OpenStudy (shaik0124):

hope this will help u

OpenStudy (anonymous):

thank you so much! would you mind helping me answer a few more math problems. Its just that I'm really stuck and I need to understand how to do this...

OpenStudy (anonymous):

wait, but thats not an answer for the multiple choice?

OpenStudy (anonymous):

Is the equation of the line \[y= \frac{ 1 }{ 3} x -3 \]

OpenStudy (shaik0124):

2nd option is answer

OpenStudy (shaik0124):

since slope is 1/3 Using the point-slope form of the equation of a line gives y - 4 = 1/3(x - 6) .

OpenStudy (shaik0124):

@Carly_Mignella

OpenStudy (shaik0124):

reply

OpenStudy (anonymous):

sorry i was on the phone with my teacher. but thank you for the help! i really appreciate it. :) it helped a lot